OpenStudy (anonymous):

4y=10 Solve this,please?

OpenStudy (anonymous):

y = 10/4 or y = 5/2 or y = 2.5

OpenStudy (anonymous):

Divide both sides by 4. 4y/4 = y and 10/4 = 2.5 so y = 2.5

OpenStudy (anonymous):